OpenStudy (anonymous):

y=2sin + sin2x when tangent is horizontal

OpenStudy (dumbcow):

tangent is horizontal when slope is 0 --> dy/dx = 0 \[2 \cos x + 2 \cos (2x) = 0\]

OpenStudy (dumbcow):

Note \[\cos (2x) = 2 \cos^2 x - 1\]